# Statistical inference on partially shape-constrained function-on-scalar linear regression models

### v1

@@ 2.4 Test procedure
We employ a resampling method to assess the statistical significance of the observed $D_{n}$ given a random sample $\mathcal{X}_{n}$. We note that resampling functional observations , say $\widetilde{\mathbf{Y}}_{i}$ $=$ $,\big{(}\widetilde{Y}_{i}(T_{i,1}),$$\ldots$,… , $\widetilde{Y}_{i}(T_{i,L_{i}})\big{)}^{\top}$, is not straightforward because $\widetilde{Y}_{i}(T_{i,1}),\ldots,\widetilde{Y}_{i}(T_{i,L_{i}})$ should inherit the joint distribution of ${Y}_{i}(T_{i,1}),\ldots,{Y}_{i}(T_{i,L_{i}})$ with within-subject heteroscedasticity. In literature, the wild bootstrap method [33, 13, 16] has been extensively investigated that can effectively handle the heteroscedasticity, and there have been many advances for handling dependent data [3, 7, 29, 11]. In this study, we use the wild bootstrap procedure as described in Algorithm 1, which can be regarded as the wild bootstrap for clustered data [3, 7].
Algorithm 1 Bootstrap procedure under the partial shape constraints
1:For a given a random sample $\mathcal{X}_{n}=\{(\mathbf{X}_{i},\mathbf{Y}_{i}):i=1,\ldots,n\}$, find the unconstrained and shape-constrained estimators $\widetilde{\boldsymbol{\beta}}$ and $\widehat{\boldsymbol{\beta}}$, respectively.
2:Set a bootstrap size $B\geq 1$.
3:Compute functional residuals;
4:for $i=1$ to $n$ do
5:     for $\ell=1$ to $L_{i}$ do
6:         $\widetilde{e}_{i,\ell}\leftarrow Y_{i,\ell}-\mathbf{X}_{i}^{\top}\widetilde{\boldsymbol{\beta}}(T_{i,\ell})$.
7:     end for
8:end for
9:Generate $B$ bootstrap samples under the null hypothesis;
10:for $b=1$ to $B$ do
11:     for $i=1$ to $n$ do
12:         for $\ell=1$ to $L_{i}$ do
13:              Sample $v_{i,\ell}^{(b)}\in\big{\{}-\frac{\sqrt{5}-1}{2},\frac{\sqrt{5}+1}{2}\big{\}}$ with probability $\Big{(}\frac{\sqrt{5}+1}{2\sqrt{5}},\frac{\sqrt{5}-1}{2\sqrt{5}}\Big{)}$.
14:              $\widetilde{Y}_{i,\ell}^{(b)}\leftarrow\mathbf{X}_{i}^{\top}\widehat{\boldsymbol{\beta}}(T_{i,\ell})+v_{i,\ell}^{(b)}\widetilde{e}_{i,\ell}$.
15:         end for
16:     end for
17:     $\mathcal{X}_{n}^{(b)}\leftarrow\{(\widetilde{\mathbf{Y}}_{i}^{(b)},\mathbf{X}_{i}):i=1,\ldots,n\}$.
18:     Find the shape-constrained $\widehat{\boldsymbol{\beta}}^{(b)}$ and unconstrained $\widetilde{\boldsymbol{\beta}}^{(b)}$ based on $\mathcal{X}_{n}^{(b)}$.
19:     Compute $D_{n}^{(b)}$ based on $\widehat{\boldsymbol{\beta}}^{(b)}$, $\widetilde{\boldsymbol{\beta}}^{(b)}$ and $\mathcal{X}_{n}^{(b)}$.
20:end for
@@
3 Simulation Study
@@
where $\mathcal{I}\subset[0,1]$ is a pre-specified sub-interval to which the partial shape constraints of interest apply. This setup is intended to mock up one of the common situations in longitudinal studies such that patients drop out of the study or transfer to other clinics. For example, as the vertical rugs show that evaluation points are right-skewed in Figure 3(b), fewer observations are available near $t=1$. Still, verifying the partial shape constraints on regression coefficient functions near $t=1$ can be of inferential interest. For a systematic assessment of the numerical experiments, we consider ten simulation scenarios of the sub-interval $\mathcal{I}=[0,0.5]$, $[0.1,0.5]$, $[0.2,0.5]$, $[0.3,0.5]$, $[0.4,0.5]$, $[0.5,1]$, $[0.6,1]$, $[0.7,1]$, $[0.8,1]$, and $[0.9,1]$. With the graphical illustration of the true regression coefficient functions in Figure 3(a), we note that the null hypothesis (3.1) is true if $\mathcal{I}\subset[0,0.5]$, and false otherwise.
@@
where $\overline{\widehat{\boldsymbol{\beta}}}(t)=R^{-1}\sum_{r=1}^{R}\widehat{\boldsymbol{\beta}}^{(r)}(t)$ is the average of the shape-constrained estimates obtained from the repeated Monte Carlo experiments. In (3.3), we normalize the ISB and IVar by $|\mathcal{I}|$ to easily compare the trend of numerical performances obtained from the different lengths of sub-intervals. Combining the above (3.2) and (3.3), we can assess the sensitivity and specificity of the proposed test procedure subject to the location and length of the sub-interval. We mainly report the simulation results obtained with $B=200$ and $R=200$, but our background simulation study gave the same lessons with larger $B$ and $R$. We refer the readers to Section S.3 of the Supplementary Materials for the implementation details, including the bandwidth and knots selection for the shape-constrained kernel smoothing and regression spline, respectively.
@@
We also consider a sub-cohort analysis for a comparative study, shedding light on another essential feature of the proposed method. As mentioned in the Introduction section, one may argue that it is more appropriate to simply treat (3.1) as global shape constraints “$\beta_{1}|_{\mathcal{I}}$ and $\beta_{3}|_{\mathcal{I}}$ are globally monotone increasing” under the model (2.1) restricted to $\mathcal{I}$. In this concern, we conduct the sub-cohort analysis using the same inferential procedure as the proposed method but only utilizing the partial data $\mathcal{X}_{n}|_{\mathcal{I}}=\{(\mathbf{Y}_{i}|_{\mathcal{I}},\mathbf{X}_{i}):i=1,\ldots,n\}$, where $\mathbf{Y}_{i}|_{\mathcal{I}}=\{Y_{i,\ell}:T_{i,\ell}\in\mathcal{I},\,\ell=1,\ldots,L_{i}\}$. In contrast to the sub-cohort analysis, we call our proposed method the “full cohort” analysis since it uses all available observations in $\mathcal{X}_{n}$ for estimation regardless of $\mathcal{I}$. As seen below, our simulation result indicates that the sub-cohort analysis suffers from a significant loss of information. The more $\mathcal{I}$ is located in the second half of the domain, the smaller the statistical power one may expect. The situation becomes much worse depending on the length of $\mathcal{I}$ and the sparsity of functional evaluations on it. Through our simulation study, we also demonstrate that the proposed test procedure is robust against the location of sub-intervals.
@@
Table 1: Simulation results for the partially shape-constrained estimates corresponding to the kernel and spline methods. The estimation and test performance are evaluated with the integrated squared bias (ISB), the integrated variance (IVar), the type I error rate, and the power of the test, defined as (3.2) and (3.3).
@@
Sample size Dataset Criterion $H_{0}:\beta_{1}$ and $\beta_{3}$ are monotone increasing on $\mathcal{I}$.
@@
$n=100$ Full data $(\gamma_{0.05},\gamma_{0.1})$ $(0.06,0.11)$ $(0.05,0.11)$ $(0.04,0.07)$ $(0.08,0.11)$
@@
Partial data $(\gamma_{0.05},\gamma_{0.1})$ $(0.06,0.11)$ $(0.09,0.19)$ $(0.05,0.11)$ $(0.06,0.12)$
restricted to $\mathcal{I}$ $\mathrm{ISB}$ $0.0517$ $0.0768$ $0.0176$ $0.1156$
@@
$n=500$ Full data $(\gamma_{0.05},\gamma_{0.1})$ $(0.05,0.10)$ $(0.05,0.09)$ $(0.04,0.08)$ $(0.05,0.08)$
@@
Partial data $(\gamma_{0.05},\gamma_{0.1})$ $(0.04,0.10)$ $(0.07,0.13)$ $(0.05,0.11)$ $(0.04,0.08)$
restricted to $\mathcal{I}$ $\mathrm{ISB}$ $0.0129$ $0.0110$ $0.0022$ $0.0109$
@@
We report the simulation results in Table 1 and Figure 4, highlighting the main lessons we obtained in the power analysis and estimation performance. The additional simulation results are provided in Section S.7 of the Supplementary Materials. The proposed partially shape-constrained inference (full cohort analysis) outperforms the sub-cohort analysis with the global shape constraints. Specifically, Table 1 shows that both the kernel and spline methods yield consistent estimates in the sense that ISB and IVar decrease as the sample size increases, which is well-aligned with the large sample properties we investigated in Section 2. Figure 4 shows that the proposed test procedure consistently meets the level of the test as specified ($\alpha=0.05,0.1$) across different scenarios varying with the length of sub-intervals from $0.1$ to $0.5$.
We note that the proposed method keeps the significance level under the null hypothesis and generally attains a greater power under the general alternative, while the sub-cohort analysis often violates the pre-specified level of the test under the null hypothesis. For the sub-interval $\mathcal{I}=[0.9,1]$, the sub-cohort analysis appears more sensitive to the general alternative than the full cohort analysis. However, as shown in Figure 1(b) in the Supplementary Materials, the shape-constrained estimates of the sub-cohort analysis suffer from significant bias due to the low sample size with $\mathcal{X}_{n}|_{\mathcal{I}=[0.9,1]}$. A similar issue arises with the sub-interval $\mathcal{I}=[0.4,0.5]$, where the sub-cohort analysis exhibits a higher type I error rate compared to the full cohort analysis. Therefore, we do not recommend using the sub-cohort analysis for verifying partial shape constraints over short sub-intervals.
Figure 4: The power analysis of the partially shape-constrained kernel method (2.9) subject to shape constraints (3.1). The sensitivity and specificity are evaluated at $5\%$ significance level under different scenarios of locations and lengths of the sub-interval as $\mathcal{I}=[0,0.5]$, $[0.1,0.5]$, $[0.2,0.5]$, $[0.3,0.5]$, $[0.4,0.5]$, $[0.5,1]$, $[0.6,1]$, $[0.7,1]$, $[0.8,1]$, and $[0.9,1]$.
@@ 4 Real Data Applications
We illustrate the application of shape-constrained inference to two datasets from clinical trials and demonstrate how we figure out treatment efficacy over time. The proposed tools help summarize dynamic efficacy patterns so that practitioners better understand when the maximum treatment effect is achieved and decide the appropriate treatment frequencies. Compared to existing studies making such conclusions relying on visualization of estimated regression coefficient functions or based on inference over the entire domain, our method enables providing statistical evidence on refined conclusions on any sub-intervals of interest. Although we illustrate examples from clinical trials, the proposed tool can be applied to any field with similar interests.
@@ 4.1 Application 1: Cervical Dystonia Dataset
Cervical dystonia is a painful neurological condition that causes the head to twist or turn to one side because of involuntary muscle contractions in the neck. Although this rare disorder can occur at any age, it most often occurs in middle-aged people, women more than men. The dataset is collected from a randomized placebo-controlled trial of botulinum toxin (botox) B, where 109 participants across 9 sites were assigned to placebo (36 subjects), 5000 units of botox B (36 subjects), or 10,000 units of botox B (37 subjects), injected into the affected muscle to partially paralyze it and make it relax. The response variable is the score on the Toronto Western Spasmodic Torticollis Rating Scale (TWSTRS), measuring the severity, pain, and disability of cervical dystonia, where higher scores mean more impairment on a 0-87 scale. The TWSTRS is administered at baseline (week 0) and at weeks 2, 4, 8, 12, and 16 thereafter. This study was originally conducted by [2], and 96.5% of subjects were followed up at designated weeks, on average. Besides, the dataset contains information on the age and sex of the subjects. The data is available in R through the package ‘medicaldata’ (https://github.com/higgi13425/medicaldata). While [2] or [8] focus on demonstrating the efficacy of botox B in cervical dystonia during weeks of clinical trial under longitudinal models, our application aims to assess the pattern of drug efficacy over weeks by applying shape-constrained inference on various sub-intervals and investigate when the maximum efficacy is achieved. To do this, we introduce the indicator variables $D_{i}$, where $D_{i}=1$ if the drug is assigned to $i$th subject, and $D_{i}=0$, otherwise. We then consider the following regression model
$Y_{i}(t)=\beta_{0}(t)+\beta_{1}(t)D_{i}+\beta_{2}(t)*\texttt{Age}_{i}+\beta_{3}(t)*\texttt{Sex}_{i}+\epsilon_{i}(t)\quad(t\in[0,16]),$ (4.1)
@@
Table 2 contains the results of the monotonicity tests, $H_{0}:\beta_{1}(t)$ is monotone decreasing (increasing) over two sub-intervals, $\mathcal{I}=[0,4]$ and $[4,16]$, from the kernel- and spline-based tests using bootstrap size $B=500$. The spline-based approach specifically adopts piece-wise quadratic $I$-spline basis functions with one internal knot located at $t=8$, determined by cross-validation as detailed in Supplementary Material S.3. To estimate $\beta_{1}(t)$ with shape restriction over $\mathcal{I}=[0,4]$ or $[4,16]$, we locate one additional knot at $t=4$ only for $\beta_{1}(t)$ to present desired shape under given boundaries. The kernel-based results are obtained from estimates based on optimal bandwidth $h\approx 3.2(=0.2\times 16)$ with the detailed selection criteria discussed in Supplementary Material. Then, we observe that both kernel- and spline-based tests conclude the significant decreasing $\beta_{1}(t)$ from weeks 0 to 4 by rejecting $H_{0}:\beta_{1}(t)$ is monotone increasing on $[0,4]$, but declare an increasing pattern observed from weeks 4 to 16 by rejecting $H_{0}:\beta_{1}(t)$ is monotone decreasing on $[4,16]$ under the significant level 0.1. This conclusion implies the improving efficacy of botox B during the first 4 weeks, but diminishing effectiveness after then. This finding becomes more convincing through consistent conclusions regardless of the choice of shape constraints on null hypotheses, either monotone decreasing or increasing, or choice of estimation methods.
@@
Figure 5 displays the estimated functional coefficients under spline estimates with the chosen optimal selection of knots in each panel. Following the inferential conclusion on treatment effect, $\beta_{1}(t)$ is estimated under the decreasing constraint over weeks from 0 to 4 and under the increasing constraint during weeks from 4 to 16. The unconstrained and constrained models are fitted using the same knots, but one extra knot at week 4 is added to fit $\beta_{1}(t)$ due to shape constraint changing at week 4. We observe the positive coefficients at week 0, presumably due to a slight bias from the random assignment in the trial, with the mean of response scores from placebo and treatment groups calculated as 43.5 and 46.7, respectively, at baseline (week 0). Although patients in the treatment group show slightly higher scores at the beginning, the decreasing trend in the first four weeks, i.e., improving efficacy, is clear from the estimated function. After 4 weeks, the degree of effectiveness weakens, and at week 16, there is no more treatment effect by returning to the status that we observed from week 0. While we find a precise fit for $\beta_{1}(t)$ under shape constraints, remainder regression coefficient estimates for age and sex, $\beta_{2}(t)$ and $\beta_{3}(t)$, show similar fits with or without constraints on $\beta_{1}(t)$. In addition to the conclusion on overall treatment efficacy as in [2], we could make a comprehensive summary of treatment efficacy based on refined conclusions over sub-intervals of interest, and it further helps precise estimation.

### v2

@@ 2.4 Testing partial shape constraints
We employ a resampling method to assess the statistical significance of the observed $D_{n}$ given a random sample $\mathcal{X}_{n}$. We note that resampling functional observations , say $\widetilde{\mathbf{Y}}_{i}$ $=$ $,\big{(}\widetilde{Y}_{i}(T_{i,1}),$$\ldots$,… , $\widetilde{Y}_{i}(T_{i,L_{i}})\big{)}^{\top}$, is not straightforward because $\widetilde{Y}_{i}(T_{i,1}),\ldots,\widetilde{Y}_{i}(T_{i,L_{i}})$ should inherit the joint distribution of ${Y}_{i}(T_{i,1}),\ldots,{Y}_{i}(T_{i,L_{i}})$ with within-subject heteroscedasticity. In literature, the wild bootstrap method [41, 15, 22] has been extensively investigated that can effectively handle the heteroscedasticity, and there have been many advances for handling dependent data [3, 7, 37, 12]. In this study, we use the wild bootstrap procedure as described in Section A of the online Supplementary Material, which can be regarded as the wild bootstrap for clustered data [3, 7].
@@
3 Numerical Study
3.1 Simulation setting
@@
where $\mathcal{I}\subset[0,1]$ is a pre-specified sub-interval to which the partial shape constraints of interest apply. This setup is intended to mock up one of the common situations in longitudinal studies such that patients drop out of the study or transfer to other clinics. For example, as the vertical rugs show that evaluation points are right-skewed in Figure 4(b), fewer observations are available near $t=1$. Still, verifying the partial shape constraints on regression coefficient functions near $t=1$ can be of inferential interest. For a systematic assessment of the numerical experiments, we consider ten simulation scenarios of the sub-interval $\mathcal{I}=[0,0.5]$, $[0.1,0.5]$, $[0.2,0.5]$, $[0.3,0.5]$, $[0.4,0.5]$, $[0.5,1]$, $[0.6,1]$, $[0.7,1]$, $[0.8,1]$, and $[0.9,1]$. With the graphical illustration of the true regression coefficient functions in Figure 4(a), we note that the null hypothesis (3.1) is true if $\mathcal{I}\subset[0,0.5]$, and false otherwise. In the numerical studies, we used an equally spaced grid of size $M=30$ on $[0,1]$ to ensure that $M_{j}\geq 3$, even in the most challenging scenario where $|\mathcal{I}_{j}|=0.1$. In our preliminary study, we also tested $M=50$ and $M=100$, and the inferential results remained consistent. Although a larger $M$ is preferable, it increases computational time. Additionally, we observed that when $M>\sum_{i=1}^{n}L_{i}$, the optimization problem often became infeasible in finite-sample studies. We recommend choosing $M$ to be as large as necessary to ensure no significant differences in (2.21), while keeping it below $\sum_{i=1}^{n}L_{i}$. In addition, we also consider the simulation studies with dummy covariates, a similar setting to the real data applications. The specific settings and results are provided in Section I of the Supplementary Material.
@@
where $\overline{\widehat{\boldsymbol{\beta}}}(t)=R^{-1}\sum_{r=1}^{R}\widehat{\boldsymbol{\beta}}^{(r)}(t)$ is the average of the shape-constrained estimates obtained from the repeated Monte Carlo experiments. In (3.3), we normalize the ISB and IVar by $|\mathcal{I}|$ to easily compare the trend of numerical performances obtained from the different lengths of sub-intervals. Combining the above (3.2) and (3.3), we can assess the sensitivity and specificity of the proposed test procedure subject to the location and length of the sub-interval. We mainly report the simulation results obtained with $B=R=250$, but our background simulation study gave the same lessons with larger $B$ and $R$. To select bandwidths and knots in a data-adaptive manner, we used a 5-fold cross-validation (CV) procedure that minimizes empirical prediction errors for individual trajectories not included in the training set. We refer the readers to Section A of the online Supplementary Material for the implementation details.
@@
For the comparison study, we first consider unconstrained estimates using data over the entire domain [0,1] but restricted to the partial domain $\mathcal{I}$. This approach is similar to the proposed method in that estimation over $\mathcal{I}$ uses the full dataset; however, the difference lies in whether the shape constraint over $\mathcal{I}$ is incorporated or not. We note that asymptotic equivalence between unconstrained and proposed partially constrained estimates is established in Theorems 2.2.2 and 2.3.1 for the kernel- and spline-based methods, respectively Then, this comparison study on estimation performance will provide insight into the finite-sample behavior of unconstrained estimates. We also consider a sub-cohort analysis for a comparative study, shedding light on another essential feature of the proposed method. As mentioned in the Introduction section, one may argue that it is more appropriate to simply treat (3.1) as global shape constraints “$\beta_{1}|_{\mathcal{I}}$ and $\beta_{3}|_{\mathcal{I}}$ are globally monotone increasing” under the model (1.1) restricted to $\mathcal{I}$. In this concern, we conduct the sub-cohort analysis using the same inferential procedure as the proposed method but only utilizing the partial data $\mathcal{X}_{n}|_{\mathcal{I}}=\{(\mathbf{Y}_{i}|_{\mathcal{I}},\mathbf{X}_{i}):i=1,\ldots,n\}$, where $\mathbf{Y}_{i}|_{\mathcal{I}}=\{Y_{i,\ell}:T_{i,\ell}\in\mathcal{I},\,\ell=1,\ldots,L_{i}\}$. In contrast to the sub-cohort analysis, we call our proposed method the “full cohort” analysis since it uses all available observations in $\mathcal{X}_{n}$ for estimation regardless of $\mathcal{I}$. As seen below, our simulation result indicates that the sub-cohort analysis suffers from a significant loss of information. The more $\mathcal{I}$ is located in the second half of the domain, the smaller the statistical power one may expect. The situation becomes much worse depending on the length of $\mathcal{I}$ and the sparsity of functional evaluations on it. Through our simulation study, we also demonstrate that the proposed test procedure is robust against the location of sub-intervals. The reproducible data generation and implementation code for kernel and spline methods are available at the following linked GitHub repository.
3.2 Simulation results
We report the simulation results in Table 1 and Figure 5, highlighting the main lessons we obtained in the power analysis and estimation performance. The additional simulation results are provided in Section H of the online Supplementary Material. First, we compare estimation performance between unconstrained and the proposed methods by evaluating ISB and IVar under a full cohort analysis with no and partial constraints. In Table 1, when $n=500$, we observe that both measures from our proposed partially constrained estimates are nearly comparable to that of the unconstrained estimates. This empirically supports the asymptotic equivalence of the two estimators, as established in Theorems 2.2.2 and 2.3.1. When $n=100$, the partially constrained estimates show a slightly larger ISB and a slightly smaller IVar compared to the unconstrained estimates, although the differences are subtle. This illustrates a potential overfitting tendency in the unconstrained estimates under finite samples, whereas the proposed method reduces variability by smoothing local variations under the given shape constraint. The slight increase in ISB for the constrained estimates aligns with the bias-variance trade-off commonly observed in nonparametric regression modeling. Next, in terms of comparison between full- and sub-cohort analyses, the proposed partially shape-constrained inference (full cohort analysis) outperforms the sub-cohort analysis with the global shape constraints. Specifically, Table 1 shows that both the kernel and spline methods yield consistent estimates in the sense that ISB and IVar decrease as the sample size increases, which is well-aligned with the large sample properties we investigated in Section 2. Figure 5 shows that the proposed test procedure consistently meets the level of the test as specified ($\alpha=0.05,0.1$) across different scenarios varying with the length of sub-intervals from $0.1$ to $0.5$.
@@ 3.2 Simulation results
Table 1: Simulation results for the partially shape-constrained estimates corresponding to the kernel and spline methods. The estimation and test performance are evaluated with the integrated squared bias (ISB), the integrated variance (IVar), the type I error rate, and the power of the test, defined as (3.2) and (3.3). In the column of Constraint, ’none’ indicates unconstrained estimation, and ‘partial’ indicates the estimation under partial shape constraint over $\mathcal{I}$.
@@
Sample size Dataset Constraint Criterion $H_{0}:\beta_{1}$ and $\beta_{3}$ are monotone increasing on $\mathcal{I}$.
@@
$n=100$ Full cohort none $\mathrm{ISB}$ $0.0288$ $0.0095$ $0.0047$ $0.0024$
$\mathrm{IVar}$ $0.4034$ $0.6507$ $0.4797$ $0.2424$
Full cohort partial $(\gamma_{0.05},\gamma_{0.1})$ $(0.06,0.11)$ $(0.05,0.11)$ $(0.04,0.07)$ $(0.08,0.11)$
@@
Sub-cohort partial $(\gamma_{0.05},\gamma_{0.1})$ $(0.06,0.11)$ $(0.09,0.19)$ $(0.05,0.11)$ $(0.06,0.12)$
subject to $\mathcal{I}$ $\mathrm{ISB}$ $0.0517$ $0.0768$ $0.0176$ $0.1156$
@@
$n=500$ Full cohort none $\mathrm{ISB}$ $0.0096$ $0.0012$ $0.0001$ $0.0001$
$\mathrm{IVar}$ $0.1067$ $0.1359$ $0.1109$ $0.0628$
Full cohort partial $(\gamma_{0.05},\gamma_{0.1})$ $(0.05,0.10)$ $(0.05,0.09)$ $(0.04,0.08)$ $(0.05,0.08)$
@@
Sub-cohort partial $(\gamma_{0.05},\gamma_{0.1})$ $(0.04,0.10)$ $(0.07,0.13)$ $(0.05,0.11)$ $(0.04,0.08)$
subject to $\mathcal{I}$ $\mathrm{ISB}$ $0.0129$ $0.0110$ $0.0022$ $0.0109$
@@
Figure 5: The power analysis of the partially shape-constrained inference with the kernel method (top) and the spline method (bottom) subject to shape constraints (3.1). The sensitivity and specificity are evaluated at $5\%$ significance level under different scenarios of locations and lengths of the sub-interval, specifically $\mathcal{I}=[0,0.5]$, $[0.1,0.5]$, $[0.2,0.5]$, $[0.3,0.5]$, and $[0.4,0.5]$ for illustration of Type I error and $\mathcal{I}=[0.5,1]$, $[0.6,1]$, $[0.7,1]$, $[0.8,1]$, and $[0.9,1]$ for illustration of power.
We note that the proposed method keeps the significance level under the null hypothesis and generally attains a greater power under the general alternative, while the sub-cohort analysis often violates the pre-specified level of the test under the null hypothesis. For the sub-interval $\mathcal{I}=[0.9,1]$, the sub-cohort analysis appears more sensitive to the general alternative than the full cohort analysis. However, as shown in Figure H.1 in the online Supplementary Material, the shape-constrained estimates of the sub-cohort analysis suffer from significant bias due to the low sample size with $\mathcal{X}_{n}|_{\mathcal{I}=[0.9,1]}$. A similar issue arises with the sub-interval $\mathcal{I}=[0.4,0.5]$, where the sub-cohort analysis exhibits a higher type I error rate compared to the full cohort analysis. Therefore, we do not recommend using the sub-cohort analysis for verifying partial shape constraints over short sub-intervals.
Our simulations, including preliminary studies with other scenarios, suggest that the kernel method has a lower Type II error rate in identifying partial shapes when the length of the restriction intervals is relatively short. We conjecture this may be partly because the spline method is slightly sensitive to the selection of knots and the number of basis functions. However, as illustrated in Figure H.3 of the Supplementary Material, we also observed that the kernel method requires significantly more computation time for large sample sizes, whereas the spline method is much faster regardless of sample size.
@@ 4 Real Data Applications
We illustrate the application of shape-constrained inference to two datasets from clinical trials and demonstrate how we figure out treatment efficacy over time. The proposed tools help summarize dynamic efficacy patterns so that practitioners better understand when the maximum treatment effect is achieved and decide the appropriate treatment frequencies. Compared to existing studies making such conclusions relying on visualization of estimated regression coefficient functions or based on inference over the entire domain, our method enables providing statistical evidence on refined conclusions on any sub-intervals of interest. Although we illustrate examples from clinical trials, the proposed tool can be applied to any field with similar interests. The reproducible code for two data examples is available at the linked GitHub repository.
@@ 4.1 Application 1: Cervical Dystonia Dataset
Cervical dystonia is a painful neurological condition that causes the head to twist or turn to one side because of involuntary muscle contractions in the neck. Although this rare disorder can occur at any age, it most often occurs in middle-aged people, women more than men. The dataset is collected from a randomized placebo-controlled trial of botulinum toxin (botox) B, where 109 participants across 9 sites were assigned to placebo (36 subjects) or botox B treatment (73 subjects), injected into the affected muscle to partially paralyze it and make it relax. The response variable is the score on the Toronto Western Spasmodic Torticollis Rating Scale (TWSTRS), measuring the severity, pain, and disability of cervical dystonia, where higher scores mean more impairment on a 0-87 scale. The TWSTRS is administered at baseline (week 0) and at weeks 2, 4, 8, 12, and 16 thereafter. This study was originally conducted by [2], and $96.5\%$ of subjects were followed up at designated weeks, on average. Besides, the dataset contains information on the age and sex of the subjects. The data is available in R through the package ‘medicaldata’ (https://github.com/higgi13425/medicaldata). While [2] or [8] focus on demonstrating the efficacy of botox B in cervical dystonia during weeks of clinical trial under longitudinal models, our application aims to assess the pattern of drug efficacy over weeks by applying shape-constrained inference on various sub-intervals and investigate when the maximum efficacy is achieved. To do this, we introduce the indicator variables $D_{i}$, where $D_{i}=1$ if botox B treatment is assigned to $i$th subject, and $D_{i}=0$, otherwise. We then consider the following regression model
$Y_{i}(t)=\beta_{0}(t)+\beta_{1}(t)D_{i}+\beta_{2}(t)*\texttt{Age}_{i}+\beta_{3}(t)*\texttt{Sex}_{i}+\epsilon_{i}(t)\quad(t\in[0,16]).$ (4.1)
@@
Table 2 contains the results of the monotonicity tests, $H_{0}:\beta_{1}(t)$ is monotone decreasing (increasing) over two sub-intervals, $\mathcal{I}=[0,4]$ and $[4,16]$, from the kernel- and spline-based tests using bootstrap size $B=500$. The spline-based approach specifically adopts piece-wise quadratic $I$-spline basis functions with one internal knot located at $t=8$, determined by cross-validation error calculation. To estimate $\beta_{1}(t)$ with shape restriction over $\mathcal{I}=[0,4]$ or $[4,16]$, we locate one additional knot at $t=4$ only for $\beta_{1}(t)$ to present desired shape under given boundaries. The kernel-based results are obtained from estimates based on optimal bandwidth $h\approx 3.2(=0.2\times 16)$ with the detailed selection criteria discussed in Section A of online Supplementary Material. Then, we observe that both kernel- and spline-based tests conclude the significant decreasing $\beta_{1}(t)$ from weeks 0 to 4 by rejecting $H_{0}:\beta_{1}(t)$ is monotone increasing on $[0,4]$, but declare an increasing pattern observed from weeks 4 to 16 by rejecting $H_{0}:\beta_{1}(t)$ is monotone decreasing on $[4,16]$ under the significant level 0.1. This conclusion implies the improving efficacy of botox B during the first 4 weeks, but diminishing effectiveness after then. This finding becomes more convincing through consistent conclusions regardless of the choice of shape constraints on null hypotheses, either monotone decreasing or increasing, or choice of estimation methods.
@@
Figure 6 displays the estimated functional coefficients under spline estimates with the chosen optimal selection of knots in each panel. Following the inferential conclusion on treatment effect, $\beta_{1}(t)$ is estimated under the decreasing constraint over weeks from 0 to 4 and under the increasing constraint during weeks from 4 to 16. The unconstrained and constrained models are fitted using the same knots, but one extra knot at week 4 is added to fit $\beta_{1}(t)$ due to shape constraint changing at week 4. We observe the positive coefficients at week 0, presumably due to a slight bias from the random assignment in the trial, with the mean of response scores from placebo and treatment groups calculated as 43.5 and 46.7, respectively, at baseline (week 0). Although patients in the treatment group show slightly higher scores at the beginning, the decreasing trend in the first four weeks, i.e., improving efficacy, is clear from the estimated function. After 4 weeks, the degree of effectiveness weakens, and at week 16, there is no more treatment effect by returning to the status that we observed from week 0. Under our proposed framework, although $p$-values from hypothesis testing quantify the degree of confidence or uncertainty about the conclusion regarding shape constraints, one may want to assess the uncertainty of the estimates. We refer to Figure R.2 in the Supplementary Material, which illustrates the 95% pointwise empirical confidence band, with calculation steps and interpretation provided there. While we find a precise fit for $\beta_{1}(t)$ under shape constraints, remainder regression coefficient estimates for age and sex, $\beta_{2}(t)$ and $\beta_{3}(t)$, show similar fits with or without constraints on $\beta_{1}(t)$. In addition to the conclusion on overall treatment efficacy as in [2], we could make a comprehensive summary of treatment efficacy based on refined conclusions over sub-intervals of interest, and it further helps precise estimation.